# “Always Nice and Confident, Sometimes wrong”: Developer’s Experiences Engaging Generative AI Chatbots Versus Human-Powered Q&A Platforms

### v2

@@ 6.2.2. Different strategies according to scenarios
Scenario 2: Problem solvingHowever, in other situations, the task quickly evolves from obtaining a simple answer to engaging in complex problem-solving. Debugging serves as a prime example; it often requires multiple iterations to pinpoint errors in the code. In such scenarios, GenAI needs to collaborate with programmers to resolve a problem, involving a more natural conversational back-and-forth process. Outlining a coding project also fits into this category. In these cases, providing detailed explanations and external materials as validity verification becomes necessary for to build trust and reach consensus. We recognize that such processes are similar to those employed in online Q&A platforms (e.g. comments, ratings, selecting correct answers, etc.). A future GenAI system could learn and emulate many strategies from platforms like SO to foster collaboration and problem-solving, rather than solely focusing on generating correct answers.
@@ 6.2.3. Combine human power with Gen AI
In our effort to compare SO with ChatGPT, we identified several specific drawbacks associated with SO. These include slower response times, a lower tolerance for questions deemed less impactful (e.g., repeated, trivial, too personalized), or those that could potentially hinder the platform’s cleanliness as a social media platform (e.g., independent follow-up questions), and the presence of a unfriendly culture. It became evident that many of these challenges could potentially be mitigated by harnessing the capabilities of AI, which may contribute to the ongoing shift in coding assistance practices. Rather than presenting future design improvements for SO, we propose a workflow that leverages the strengths of both platforms.
SO, functioning as a blend of Q&A and social media, differs in its operation from ChatGPT. We acknowledge the potential disparities in usage phases between these two platforms. ChatGPT, being an ‘always-on’ personal assistant, serves as the initial phase of assistance. Developers could get fast and clear answers without too much effort. Upon receiving answers from AI after several iterations, in cases where the problem remains unsolved or users harbor doubts about response accuracy, we propose introducing external validation and assistance. In Phase II, developers seek crowd guidance by directly reporting issues to Q&A platforms like SO. Other software engineers can then participate in the decision-making process by voting to verify answer accuracy, providing comments, and offering additional answers to both the initial and follow-up questions. Given that SO banned AI-generated answers partly due to identification difficulties, we suggest introduce specific tags like ”AI-generated”, for posts to enhance clarity and categorization.
@@
In this workflow, AI chatbot handles tedious or too personalized tasks, while Q&A serves as an external assistance that maintains its own affordances as social platforms. While this workflow is by no means perfect, we aspire to offer inspiration on how future systems could potentially harness the strengths of both platforms to further enhance developer assistance in the coding process.
@@ 6.3. From the past to the future: The evolving role of GenAI in coding assistance
In the previous section, we presented a sample workflow for seeking programming guidance through both AI-powered chatbots and human-powered Q&A platforms. We aimed to leverage the strengths of both platforms, but we also pondered whether, with researchers continually enhancing AI to harness human creativity and tackle complex tasks, it might eventually replace platforms like SO and other Q&A sites. While it’s impossible to predict the future with certainty, we do believe that AI will play a significant role in replacing some human support. However, it’s crucial to note that current AI systems still learn from human-generated data, even for GenAI that can create content. As demonstrated by Xue et al., the lack of user sharing open-sourced information could pose challenges to the sustainability of LLM’s learning process (Xue et al., 2023). Therefore, we suggest that future systems strike a balance, not only for the sustainable improvement of AI but also to preserve the social aspects tied to Q&A platforms. These platforms should continue to be spaces where users not only seek answers but also engage in a community to collectively find solutions.
@@ 6.4. Limitations
While we aimed to ensure the robustness of each research step, there are still some limitations to this work. Although ChatGPT has garnered considerable attention during the specified timeframe, comprehensively capturing individuals’ complete experiences with it through Reddit posts analysis remains challenging due to factors such as the silent effect and potential biases (dar, 2012). Our data analysis strategy has been carefully devised to accurately capture the prevailing trends, but augmenting our resources could lend greater robustness to our results. In the future, conducting follow-up comparative user studies would be valuable to further validate our insights.
With enhanced resources, addressing the technical intricacies could involve refining the search process by extending it beyond the posts’ titles and selftext to encompass the entire content, including comments. Moreover, the main changes of Reddit APIs and the shut down of pushshift.io impacted our ability to access more expansive datasets. Readers should also be aware that the accuracy of the API is not guaranteed to be 100% primarily due to potential delays (pra, 2023). Nonetheless, this approach remains the most accurate within the scope of research (Gauthier et al., 2022). It’s worth noting that our results have exhibited a greater degree of promise since our emphasis toward qualitative analysis.
@@ 7. Conclusion
Appendix A Appendix

### v3

@@ 6.2.2. Different Strategies According to Scenarios
Scenario 2: Problem-solvingHowever, in other situations, the task quickly evolves from obtaining a simple answer to engaging in complex problem-solving. Debugging serves as a prime example; it often requires multiple iterations to pinpoint errors in the code. In such scenarios, GenAI needs to collaborate with programmers to resolve a problem, involving a more natural conversational back-and-forth process. Outlining a coding project also fits into this category. In these cases, providing detailed explanations and external materials as validity verification becomes necessary to build trust and reach consensus. We recognize that such processes are similar to those employed in online Q&A platforms (e.g. comments, ratings, selecting correct answers, etc.). A future GenAI system could learn and emulate many strategies from platforms like SO to foster collaboration and problem-solving, rather than solely focusing on generating correct answers.
@@ 6.2.3. Combine Human Power with Gen AI
In our effort to compare SO with ChatGPT, we identified several specific drawbacks associated with SO. These include slower response times, a lower tolerance for questions deemed less impactful (e.g., repeated, trivial, too personalized), or those that could potentially hinder the platform’s cleanliness as a social media platform (e.g., independent follow-up questions), and the presence of an unfriendly culture. It became evident that many of these challenges could potentially be mitigated by harnessing the capabilities of AI, which may contribute to the ongoing shift in coding assistance practices. Rather than presenting future design improvements for SO, we propose a workflow that leverages the strengths of both platforms.
SO, functioning as a blend of Q&A and social media, differs in its operation from ChatGPT. We acknowledge the potential disparities in usage phases between these two platforms. ChatGPT, being an ‘always-on’ personal assistant, serves as the initial phase of assistance. Developers could get fast and clear answers without too much effort. Upon receiving answers from AI after several iterations, in cases where the problem remains unsolved or users harbor doubts about response accuracy, we propose introducing external validation and assistance. In Phase II, developers seek crowd guidance by directly reporting issues to Q&A platforms like SO. Other software engineers can then participate in the decision-making process by voting to verify answer accuracy, providing comments, and offering additional answers to both the initial and follow-up questions. Given that SO banned AI-generated answers partly due to identification difficulties, we suggest introducing specific tags like ”AI-generated”, for posts to enhance clarity and categorization.
@@
In this workflow, AI chatbot handles tedious or too personalized tasks, while Q&A serves as an external assistant that maintains its own affordances as a social platform. While this workflow is by no means perfect, we aspire to offer inspiration on how future systems could potentially harness the strengths of both platforms to further enhance developer assistance in the coding process.
@@ 6.3. From the Past to the Future: The Evolving Role of GenAI in Coding Assistance
In the previous section, we presented a sample workflow for seeking programming guidance through both AI-powered chatbots and human-powered Q&A platforms. We aimed to leverage the strengths of both platforms, but we also pondered whether, with researchers continually enhancing AI to harness human creativity and tackle complex tasks, it might eventually replace platforms like SO and other Q&A sites. While it’s impossible to predict the future with certainty, we do believe that AI will play a significant role in replacing some human support. However, it’s crucial to note that current AI systems still learn from human-generated data, even for GenAI which can create content. As demonstrated by Xue et al., the lack of user sharing open-sourced information could pose challenges to the sustainability of LLM’s learning process (Xue et al., 2023). Therefore, we suggest that future systems strike a balance, not only for the sustainable improvement of AI but also to preserve the social aspects tied to Q&A platforms. These platforms should continue to be spaces where users not only seek answers but also engage in a community to collectively find solutions.
@@ 6.4. Limitations
While we aimed to ensure the robustness of each research step, there are still some limitations to this work. Although ChatGPT has garnered considerable attention during the specified timeframe, comprehensively capturing individuals’ complete experiences with it through Reddit post analysis remains challenging due to factors such as the silent effect and potential biases (dar, 2012). Our data analysis strategy has been carefully devised to accurately capture the prevailing trends, but augmenting our resources could lend greater robustness to our results. In the future, conducting follow-up comparative user studies would be valuable to further validate our insights.
With enhanced resources, addressing the technical intricacies could involve refining the search process by extending it beyond the posts’ titles and selftext to encompass the entire content, including comments. Moreover, the main changes in Reddit APIs and the shutdown of pushshift.io impacted our ability to access more expansive datasets. Readers should also be aware that the accuracy of the API is not guaranteed to be 100% primarily due to potential delays (pra, 2023). Nonetheless, this approach remains the most accurate within the scope of research (Gauthier et al., 2022). It’s worth noting that our results have exhibited a greater degree of promise since our emphasis on qualitative analysis.
@@ 7. Conclusion
8. Acknowledgment
We are grateful to our collaborators for all their help and support throughout this project. Their knowledge and guidance shaped our research and improved its overall impact. We also appreciate the reviewers’ careful reading and thoughtful comments, which prompted us to clarify our ideas and present our findings more effectively.
Appendix A Codebooks with Themes and Quote Examples